Assume that z is the test statistic.

(a) H0: μ = 22.5, Ha: μ > 22.5; x = 25.9, σ = 7.4, n = 33
(i) Calculate the test statistic z. (Round your answer to two decimal places.)


(ii) Calculate the p-value. (Round your answer to four decimal places.)


(b) H0: μ = 200, Ha: μ < 200; x = 193.8, σ = 35, n = 36
(i) Calculate the test statistic z. (Round your answer to two decimal places.)


(ii) Calculate the p-value. (Round your answer to four decimal places.)


(c) H0: μ = 12.4, Ha: μ ≠ 12.4; x = 11.3, σ = 5.3, n = 26
(i) Calculate the test statistic z. (Round your answer to two decimal places.)


(ii) Calculate the p-value. (Round your answer to four decimal places.)
Answer #1

(a) H0: μ = 22.5, Ha: μ > 22.5;    This is right tailed test

Given that, x = 25.9,        σ = 7.4,            n = 33

(i) Therefore, the test statistic for the problem will be,

r_μ 25.9-22.5 3.4 σ/Vn 7.4/V33 1.2882 = 2.6394
(ii) The p-value for the right tailed test corresponds to probability P(Z > z) which can be obtained using Excel function NORMSDIST(z) as below:
p - value-P(Z > 21- P(Z)1- NORMSDIST (2.6394) (1- 0.9958) 0.0042

(b) H0: μ = 200, Ha: μ < 200;    This is left tailed test

Given that, x = 193.8,        σ = 35,            n = 36

(i) Therefore, the test statistic for the problem will be,

.-2006.1.0629 x-11 193.8-200 =-1.0629 = a/vn35/36583331.0629 5 8333

(ii) The p-value for the left tailed test corresponds to probability P(Z 2) which can be obtained using Excel function NORMSDIST(z) as below:
p-value = P(Z < z) = NORMSDIST(-1.0629) = 0.1439

(c) H0: μ = 12.4, Ha: μ ≠ 12.4    This is left tailed test

Given that, x = 11.3,        σ = 5.3,            n = 26

(i) Therefore, the test statistic for the problem will be,

_ μ 11.3-12.4 一一1.0583 o/vm 5.3/V26 = 1.0394

(ii) The p-value for the two tailed test corresponds to probability 2P(Z leq z) which can be obtained using Excel function NORMSDIST(z) as below:
p-value=2P(Zleq z) =2NORMSDIST(-1.0583)=2(0.1450)=0.2899